Solve for x
Type an integer or simplified fraction, and use a comma to separate answers.
6x^2 = -25x - 4
Thanks very much!!

Answer by scott8148(6628)

6x^2 + 25x + 4 = 0

factoring ___ (6x + 1)(x + 4) = 0

6x + 1 = 0 ___ 6x = -1 ___ x = -1/6

x + 4 = 0 ___ x = -4